A reaction is followed and found to have a rate constant of 3.36x10^4 M^-1 s^-1 at 344K and a rate constant of 7.69 M^-1 s^-1 at 219K. Determine the activation energy for this reaction.


1
Expert's answer
2021-03-10T18:27:17-0500

"k=A e^{-E\/(RT)}"

"k_1\/k_2=e^{E\/(RT_2)-E\/(RT_1)}"

"ln \left (\frac {k_1} {k_2}\right) =\frac {ET_1 -ET_2} {RT_1T_2}"

"E=\frac{RT_1T_2} {T_1-T_2}ln\left(k_1/k_2\right)"

"E=42 kJ\u2022mol^{-1}"